OpenStudy (anonymous):

Can someone help me simplify (2 square root 5 + 3 square root 7)^2

OpenStudy (anonymous):

\[(2\sqrt{5}+3\sqrt{7})(2\sqrt{5}+3\sqrt{7})\] and now 4 multiplications

OpenStudy (anonymous):

What do you mean by 4 multiplications?

OpenStudy (anonymous):

\[2\sqrt{5}\times 2\sqrt{5}+2\times 2\sqrt{5}\times 3\sqrt{7}+3\sqrt{7}\times 3\sqrt{7}\] \[4\times 5+12\sqrt{35}+9\times 7\] \[20+12\sqrt{35}+63\] \[83+12\sqrt{35}\]

OpenStudy (anonymous):

\[(a+b)^2=a^2+ab+ab+b^2=a^2+2ab+b^2\] is what i meant

OpenStudy (anonymous):

OHH okay, got it. so then whats after that?

OpenStudy (anonymous):

done, i wrote all the steps in all the details above

OpenStudy (anonymous):

oh thats it? wow okay thanks again for your help :)